OpenStudy (anonymous):

What is the next term for the given arithmetic sequence? -3, -2.25, -1.5, -0.75, ... -0.25 0 0.75 0??

OpenStudy (anonymous):

find the common difference

OpenStudy (akashdeepdeb):

Okay so in this arithmetic sequence The first term is -3 And the common difference is?

OpenStudy (anonymous):

it moves 0.75

OpenStudy (anonymous):

yeah +0.75

OpenStudy (akashdeepdeb):

Correct so how much would it be after it moved +0.75 aftre - 0.75??

OpenStudy (anonymous):

-3 + 0.75 = -2.25 -2.25 + 0.75 = -1.50 -1.50 + 0.75 = -0.75 -0.75 + 0.75 = ?

OpenStudy (akashdeepdeb):

Yes you are right it IS 0!!

OpenStudy (anonymous):

ex: -1 + 1 = 0

OpenStudy (anonymous):

medal?